Suppose Quantity Demanded is given by Qd(p) = 25 - 2 x p and Quantity Supplied...

  • Suppose Quantity Demanded is given by Qd(p) = 25 - 2 x p and Quantity Supplied is given by Qs(p) = 11 + 5 x p If the price is $3.00
  • Answer = There will be excess supply (a surplus)
  • can you please explain the math behind this? how to calculate for surplus?
Answer #1

Suoplus = Supply demand = 11+ 5p - (25-20) Wilhen p= 3 - 11 + 5p -25 + 2p 7-14 When P= 3 7x3-14 21-14 - Suoplus = 7 units